OpenStudy (anonymous):

m=4/5 and (1,-6) the equation of the line is y=4/5x - ?/5

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

i dont know. can you help me? know how to calculate it. can you help me?

jimthompson5910 (jim_thompson5910):

y = mx + b y = 4/5x + b ... plug in the given slope m = 4/5 -6 = (4/5)(1) + b ... plug in x = 1 and y = -6 -6 = 4/5 + b solve for b and tell me what you get

OpenStudy (anonymous):

-34/5

jimthompson5910 (jim_thompson5910):

good